OpenStudy (anonymous):

Someone please helpppIs L(x,y)=(x+y,x-y) a linear transformation?(Show with full working)

OpenStudy (perl):

if L(u + v) = L(u) + L(v) , and L ( c*u) = c* L(u), then its a linear transformation

OpenStudy (perl):

and the definition says a function from R^n to R^m, here n=m = 2

OpenStudy (anonymous):

Yup i know the conditions but i dont know how to apply it.

OpenStudy (anonymous):

Could you help me solve in terms of two vectors. For both right side and left side

OpenStudy (perl):

L(x,y)=(x+y,x-y) Let u = (u1, u2) v = (v1,v2) Then u + v = ( u1 + v1 , u2 + v2) Therefore L ( u + v ) = L ( u1+v1, u2 + v2) = ( u1 + v1 + u2 + v2 , u1 + v1 - ( u2 + v2) )

OpenStudy (anonymous):

okay is this the final answer ?? or we have to check scalar multiplication too?

OpenStudy (perl):

now its not clear how to proceed, lets start from the other side evaluate L(u) + L(v)

OpenStudy (anonymous):

yes please im watching

OpenStudy (anonymous):

u there??

OpenStudy (perl):

u = (u1,u2) v = (v1,v2) L( u ) = ( u1 + u2, u1 - u2 ) L( v) = (v1 + v2 , v1 - v2) L(u) + L(v) = ( u1 + u2, u1 - u2 ) + (v1 + v2 , v1 - v2) = ( u1 + u2 + v1 + v2 , u1 - u2 + v1 - v2) = ( u1 + v1 + u2 + v2, u1 + v1 - (u2 + v2) )

OpenStudy (perl):

now the last expression, that is what we got for L ( u + v )

OpenStudy (anonymous):

okayy...now its much clear

OpenStudy (anonymous):

do we need to chk multiplication also

OpenStudy (anonymous):

thanks mate

OpenStudy (perl):

hey, how about this

OpenStudy (perl):

you know vectors are bold, but vector components are normal shaded

OpenStudy (anonymous):

yup

OpenStudy (perl):

i cant draw bold vectors, so i will use capital letters

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

Based on perl's definitions, we can say L ( u + v ) = L ( u1+v1, u2 + v2) L ( u + v ) = ( u1 + v1 + u2 + v2 , u1 + v1 - ( u2 + v2) ) L ( u + v ) = ( u1 + v1 + u2 + v2 , u1 + v1 - u2 - v2) ) L ( u + v ) = ( (u1 + u2) + (v1 + v2) , (u1 - u2) + (v1 - v2)) ) L ( u + v ) = (u1 + u2,u1 - u2) + (v1 + v2,v1 - v2) L ( u + v ) = L(u) + L(v) So that takes care of the first part. You need to do the same for L(c*u) = c*L(u)

OpenStudy (perl):

Given L(x,y)=(x+y,x-y) show L is a linear transformation from R^2 to R^2 Let U = (u1, u2) V = (v1,v2) Then U + V = ( u1 + v1 , u2 + v2) Therefore L(U) + L(V) = ( u1 + u2, u1 - u2 ) + (v1 + v2 , v1 - v2) = ( u1 + u2 + v1 + v2 , u1 - u2 + v1 - v2) = ( u1 + v1 + u2 + v2, u1 + v1 - (u2 + v2) ) = L(u1 + v1, u2 + v2) = L ( U + V )

OpenStudy (anonymous):

ohhhhhhhhhhhh i get that thanks jim n perl. you rock

OpenStudy (perl):

L( c*U) = ( c*u1 + c*u2 , c*u1 - c*u2) = c * ( u1 + u2 , u1 - u2) = c* L( U )

OpenStudy (anonymous):

thanks a million